charges of 3nC, -2nC, -7nC and 1nC are contained inside a rectangular box with length 1m width 2m and height 2.5m. Outside the boc are charges of 1nC and 4nC. What os the electric flux through the sirface.
1
Expert's answer
2019-10-09T10:46:55-0400

The total of the electric flux for a closed surface is equal to the total charge divided by the permittivity.

In our case the total charge Q=Qinside-Qoutside=(3nC-2nC-7nC+1nC)-(1nC+4nC)=-10nC

Total flux F=Q/e0=-10nC/(8.85*10-12F/m)=-1.13*103V*m